### v1

@@ Abstract
Population synthesis simulations of compact binary coalescences (CBCs) play a crucial role in extracting astrophysical insights from an ensemble of gravitational wave (GW) observations. However, realistic simulations can be costly to implement for a dense grid of initial conditions. Normalizing flows can emulate the distribution functions of a simulated population of binary parameters and thereby enable empirical constraints on the astrophysical initial conditions and branching fractions of various formation channels, given data from a catalog of GW observations. They can also be used for data amplification in sparse regions of the CBC parameter space to better guide the development of phenomenological population models for rarely synthesizable systems with components in theorized mass gaps, without having to simulate a prohibitively large number of binaries. However, flow predictions are wrought with uncertainties, especially for sparse training sets. In this work, I develop a method for quantifying and marginalizing uncertainties in the emulators by introducing the Bayesian Normalizing flow, a conditional density estimator constructed from Bayesian neural networks. Using the exact likelihood function naturally associated with density estimators, I sample the posterior distribution of flow parameters with suitably chosen priors to quantify and marginalize over flow uncertainties. I demonstrate the accuracy, calibration, and data-amplification impacts of the estimated uncertainties for simulations of binary black hole populations formed through common envelope evolution. I outline the applications of the proposed methodology in the context of simulation-based inference from growing GW catalogs and feature prediction, with state-of-the-art binary evolution simulators, now marginalized over model and data uncertainties. I also sketch other applications for general simulation-based approaches in GW astronomy.
@@ 1 Introduction
Population synthesis (PopSynth, K. Breivik, 2025) simulations can model these astrophysical imprints on the various features and correlations of the population-level distributions of measurable CBC parameters. PopSynth predictions can guide the construction of targeted phenomenological models (e.g., C. Talbot & E. Thrane, 2018; L. A. C. van Son et al., 2022) of the distribution function whose parameters can be constrained given GW data from growing catalogs to extract astrophysical insights (R. Abbott et al., 2023a). More flexible population models (e.g., B. Edelman et al., 2023; T. A. Callister & W. M. Farr, 2024; A. Ray et al., 2023, 2024; J. Heinzel et al., 2025) are also employed to search for additional physics in the data beyond the assumptions of astrophysically motivated density functions, with the reconstructed population features being harder to interpret in terms of the physical processes underlying an evolutionary channel. On the other hand, direct comparison of PopSynth outputs to a collection of measured CBC parameters by means of density estimation can provide observational constraints on astrophysical initial conditions of the simulated evolutionary pathways (M. Zevin et al., 2021; J. Riley & I. Mandel, 2023; S. Colloms et al., 2025). Such approaches can also be considered reasonably flexible in terms of the underlying physical scenarios explored provided a diverse collection of simulated formation channels are used in the inference.
@@
Normalizing flows (NFs, D. J. Rezende & S. Mohamed, 2016; G. Papamakarios et al., 2021) are powerful generative models that can assist with both simulation-based inference (SBI) and feature prediction, from PopSynth outputs. For example, S. Colloms et al. (2025) use NFs for estimating the conditional distribution of CBC parameters for each simulated population given values of the astrophysical initial conditions that characterize the simulated pathways. Once trained, a weighted mixture of different estimators, each corresponding to a specific formation channel, can facilitate tractable interpolation of the population likelihood across discrete gridpoints of simulation initial conditions, which can be constrained simultaneously with the branching fractions (mixture weights) of each channel given GW data from growing catalogs. S. Colloms et al. (2025) constrain the common envelop ejection effeciency, an uncertain parameter in isolated binary stellar evolution, the natal spin of black holes (BHs), and the branching fractions of all five formation channels whose simulations are used in their inference, from the confident binary black hole (BBH) detections (excluding population outliers) reported in the third GW transient catalog (GWTC-3 R. Abbott et al., 2023b). On the other hand, NF emulators can also be used for data amplification to predict population-level features and correlations in underrepresented regions of the CBC parameter space, where very few systems are synthesized from a particular channel, without having to simulate a prohibitively large ensemble of binaries. However, like other deep learning emulators, the predictions of normalizing flows are susceptible to epistemic (model) and aleatoric (data) uncertainties (V. Nemani et al., 2023; W. He et al., 2025), which can, in principle, bias the inferred astrophysical conclusions.
@@
For flow-based emulators, model uncertainties may arise either from inefficiencies in parameter learning, sub-optimal architectures, and out of distribution samples in the observed data, where as data uncertainties may result from the inherent sparsity and randomness of the training set which may be caused by particular formation channels being highly inefficient in certain regions of the CBC parameter space or by numerical approximation used by the simulators internally to speed up computation (V. Nemani et al., 2023; W. He et al., 2025; J. J. Andrews et al., 2024). While biases associated with ignoring model uncertainties may, in principle, be reduced by generating more training data, costly simulators can prohibit such endeavors. On the other hand, astrophysical bottlenecks leading to sparse and noisy datasets cannot be mitigated by simulating a larger number of binaries. In other words, for state of the art (costly) simulators and extended investigations that include observations of rarely synthesizable systems such as highly unequal mass neutron Star black hole binaries (NSBHs)/BBHs with a component in the $3-5M_{\odot}$ mass range (R. Abbott et al., 2020; M. Zevin et al., 2020; A. G. Abac et al., 2024), it is necessary to quantify and marginalize over emulator uncertainties for ensuring unbiased astrophysical inference.
@@
In this paper, I introduce the Bayesian normalizing flow, a framework for emulating conditional and unconditional distribution functions equipped with robust uncertainty quantification and model comparison among architectures. By constructing the transformation layers of an NF from Bayesian neural networks (BNNs) and exploiting the natural likelihood function associated with any density estimator model, such as an NF, I sample the posterior distribution of flow parameters given training data and suitable a priori assumptions. Using these samples, I predict the emulated population as a Bayesian credible interval, rather than a deterministic density function, which is expected to contain the true population-level distribution of synthesized CBC parameters with a certain posterior probability. Using population synthesis simulations of BBHs formed through common envelope (CE) evolution of isolated stellar binaries (S. S. Bavera et al., 2020), I show that the inferred credible intervals correctly quantify epistemic and aleatoric uncertainties in flow predictions by computing calibration curves, and by also comparing the posterior-predictive distributions with those of multiple normal training re-runs on alternate realizations of sparse (thinned) datasets or on the same dataset with different parameter initializations. I further show how to compute the Bayesian information criterion for different flow architectures to enable model comparison and thereby reduce the epistemic uncertainties that arise from sub-optimal architecture choices.
In a previous work, S. Bieringer et al. (2024) attempted Bayesian uncertainty quantification (UQ) for unsupervised (unconditional) density estimation using continuous normalizing flows (C. Chen et al., 2018; W. Grathwohl et al., 2018). They have either used approximate loss functions (Y. Lipman et al., 2023) instead of the exact log-likelihood or highly fine-tuned sampling methods (S. Bieringer et al., 2023) that have been subject to limited testing, to conduct their inference. They further restrict to a few (O(50)) posterior samples for constructing calibration statistics and do not implement any model-comparison framework. Instead, the method proposed here uses more robust, stable and rigorously tested sampling techniques based on Hamiltonian Monte Carlo (HMC, M. Betancourt, 2018; M. D. Hoffman & A. Gelman, 2011) methods, exact log-likelihoods, and calibration statistics constructed from thousands of posterior samples, which ensure stability and convergence. While I use masked autoregressive flows (G. Papamakarios et al., 2018) for the proof-of-concept demonstration presented here, this method is easily generalizable to continuous flows, which are expected to enable much faster convergence for HMC sampling. As I show later on, for advanced continuous flow algorithms (such as FFJORD, W. Grathwohl et al., 2018) with scalable log density evaluations, the complexity required to emulate realistic CBC PopSynth simulations is also well within the scope of modern HMC samplers. Hence, a PopSynth emulator based on Bayesian continuous flows would not necessitate highly fine-tuned samplers such as the ones proposed by S. Bieringer et al. (2023). These endeavors, along with other improvements and generalizations, such as applications to other simulation-based approaches in GW astronomy (M. Dax et al., 2021, 2025), are left as future explorations.
@@
The rest of the paper is organized as follows. In Sec. 2, I highlight the main motivations for developing this framework. In Sec. 4, I describe in detail the Bayesian normalizing flow, its training, implementation, and calibration. In Sec. 5, I demonstrate the accuracy and interpretability of the quantified uncertainties for a fiducial BBH PopSynth dataset while also highlighting the data amplification potential of PopSynth emulators. I conclude in Sec. 6 with a summary of the main developments presented and the follow-up investigations planned for the near future.
2 Motivation: the necessity of UQ for state-of-the-art simulators
@@
The scope and reliability of SBI and feature prediction are determined by the fidelity of the simulator and its underlying physical prescriptions. Compared to its earlier versions (S. S. Bavera et al., 2020), which were used to model some isolated BBH formation scenarios in the training sets of S. Colloms et al. (2025), the POSYDON framework (T. Fragos et al., 2023; J. J. Andrews et al., 2024) introduced significant advances in the treatment of binary stellar evolution, greatly expanding the astrophysical scope of SBI from GWs. POSYDON provides one of the most accurate descriptions of isolated CBC formation available in the literature by relying on extensive grids of binary evolution models that track stellar structure for both components using MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019) from zero-age main sequence to compact object formation, combined with advanced interpolation methods (T. Fragos et al., 2023). In POSSYDONv2, further improvements are achieved by accounting for evolution across a cosmological range of metallicities, reverse mass transfer, spin-orbit misalignment and binary disruption due to supernova kicks, and the possibility of stellar mergers (J. J. Andrews et al., 2024), etc. A significant improvement in scalability, while retaining accuracy, is also achieved through the implementation of efficient machine learning algorithms. Altogether, POSYDONv2 is emerging as a tracktable PopSynth simulator for isolated CBC formation (and many other astrophysical systems of interest) which boasts one of the highest physics fidelities to date.
@@
Therefore, given the high computational cost of running a detailed binary simulation, POSYDONv2 can be assisted by NF emulators not only for SBI but also for feature prediction. For the latter, NFs can be used to amplify the data, which amounts to boosting sample statistics for a particular population, by learning from many smaller simulation sets at nearby initial condition grid points. In other words, instead of having to simulate millions of binaries for a single intitial condition of interest to properly sample under-represented regions of the CBC parameter space, many small simulation sets on a grid of nearby initial condition values (that are paralelizable as individual POSYDON runs), can be used to train an NF emulator, which can then generate a large reliable sample of binaries for the particular initial conditions of interest. This can facilitate scalable and accurate population feature-prediction for rarely synthesizable CBCs without having to simulate a prohibitively large number of systems. I show this later on for an existing dataset simulated with earlier versions of POSYDON, as proof of concept.
@@
To summarize, PopSynth emulators trained on a diverse collection of POSYDONv2 runs promise novel and reliable astrophysical insights from growing CBC catalogs. However, generating arbitrarily many simulated binaries per initial condition grid point and using them to train emulators are both prohibitive, given the high computational cost of a single PopSynth run and that of NFs attempting to learn from billions or more datapoints, respectively. The latter might necessitate further downsampling of CBCs per population run in the SBI and/or feature-prediction training set. While NFs can facilitate data amplification for a particular simulation grid point by learning from others (as I shall show later on), the predicted populations and density evaluations will be uncertain upon re-training with alternate realizations of down-sampled data or alternate initializations of flow parameters. The inherent randomness in POSYDON outputs due to numerical interpolation schemes used internally to speed up population synthesis, and sparsity in certain regions of CBC parameter space due to astrophysical bottlenecks, can be expected to further contribute to the variance in flow predictions for downsampled datasets. Emulating POSYDON runs for SBI and feature predictions, therefore, necessitates robust uncertainty quantification and marginalization to ensure reliable astrophysical explorations.
3 Methods

### v2

@@ Abstract
Population synthesis simulations of compact binary coalescences (CBCs) play a crucial role in extracting astrophysical insights from an ensemble of gravitational wave (GW) observations. However, realistic simulations can be costly to implement for a dense grid of initial conditions. Normalizing flows can emulate population synthesis runs to enable simulation-based inference from observed catalogs and data augmentation for feature prediction in rarely synthesizable sub-populations. However, flow predictions can be wrought with uncertainties, especially for sparse training sets. In this work, we develop a method for quantifying and marginalizing uncertainties in the emulators by implementing the Bayesian Normalizing flow, a conditional density estimator constructed from Bayesian neural networks. Using the exact likelihood function naturally associated with density estimators, we sample the posterior distribution of flow parameters with suitably chosen priors to quantify and marginalize over flow uncertainties. We demonstrate the accuracy, calibration, inference, and data-augmentation impacts of the estimated uncertainties for simulations of binary black hole populations formed through common envelope evolution. We outline the applications of the proposed methodology in the context of simulation-based inference from growing GW catalogs and feature prediction, with state-of-the-art binary evolution simulators, now marginalized over model and data uncertainties.
@@ 1 Introduction
Population synthesis (PopSynth, see K. Breivik, 2025, for a review) simulations can model these astrophysical imprints on the various features and correlations of the population-level distributions of measurable CBC parameters. PopSynth predictions can guide the construction of targeted phenomenological models (e.g., C. Talbot & E. Thrane, 2018; L. A. C. van Son et al., 2022) of the distribution function whose parameters can be constrained given GW data from growing catalogs to extract astrophysical insights (R. Abbott et al., 2023). More flexible population models (e.g., B. Edelman et al., 2023; T. A. Callister & W. M. Farr, 2024; A. Ray et al., 2023, 2024; J. Heinzel et al., 2025) are also employed to search for additional physics in the data beyond the assumptions of astrophysically motivated density functions, with the reconstructed population features being harder to interpret in terms of the physical processes underlying an evolutionary channel. On the other hand, direct comparison of PopSynth outputs to a collection of measured CBC parameters by means of density estimation can provide observational constraints on astrophysical initial conditions of the simulated evolutionary pathways (M. Zevin et al., 2021; A. Q. Cheng et al., 2023; J. Riley & I. Mandel, 2023; S. Colloms et al., 2025; S. Mastrogiovanni et al., 2022; Y. Bouffanais et al., 2019, 2021; K. W. K. Wong et al., 2021; M. Mould et al., 2022; C. Plunkett et al., 2025). Such approaches can also be considered reasonably flexible in terms of the underlying physical scenarios explored provided a diverse collection of simulated formation channels are used in the inference.
@@
Normalizing flows (NFs, D. J. Rezende & S. Mohamed, 2016; G. Papamakarios et al., 2021) are powerful generative models that can assist with both simulation-based inference (SBI) and feature prediction, from PopSynth outputs. For example, S. Colloms et al. (2025); C. Plunkett et al. (2025) use NFs for estimating the conditional distribution of CBC parameters for each synthetic population given values of the astrophysical initial conditions that characterize the simulated pathways. They perform simulation-based inference (SBI) by efficiently interpolating the population likelihood over a grid of astrophysical initial conditions to constrain their posterior distribution given observational data from GW catalogs. On the other hand, NF emulators can also be used for data augmentation to predict population-level features and correlations in underrepresented regions of the CBC parameter space, where very few systems are synthesized from a particular channel, without having to simulate a prohibitively large ensemble of binaries (as we demonstrate later on). However, like other deep learning emulators, the predictions of normalizing flows are susceptible to epistemic (model) and aleatoric (data) uncertainties (V. Nemani et al., 2023; W. He et al., 2025), which can, in principle, bias the inferred astrophysical conclusions.
@@
For flow-based emulators, model uncertainties may arise either from inefficiencies in parameter learning, sub-optimal architectures, and out of distribution samples in the observed data, where as data uncertainties may result from the inherent sparsity and randomness of the training set which may be caused by particular formation channels being highly inefficient in certain regions of the CBC parameter space or by numerical approximation used by the simulators internally to speed up computation (V. Nemani et al., 2023; W. He et al., 2025; J. J. Andrews et al., 2024). While biases associated with ignoring model uncertainties may, in principle, be reduced by generating more training data, costly simulators can prohibit such endeavors. On the other hand, astrophysical bottlenecks leading to sparse and noisy datasets are difficult to mitigate against whilst simulating a tractable number of binaries. In other words, for SBI with state-of-the-art (costly) simulators or to correctly characterize their theoretical predictions for rarely synthesizable sub-populations such as highly unequal mass neutron Star black hole binaries (NSBHs), BBHs with a component in the $3-5M_{\odot}$ mass range (R. Abbott et al., 2020; M. Zevin et al., 2020; A. G. Abac et al., 2024), CBCs emerging from stellar binaries at very high metallicities (K. Belczynski et al., 2010), and synthetic populations corresponding to large natal kicks etc., it is necessary to quantify and marginalize over emulator uncertainties for unbiased astrophysical inference.
In this paper, we implement the Bayesian normalizing flow, a framework for emulating conditional and unconditional distribution functions equipped with uncertainty quantification and model comparison among architectures. By constructing the transformation layers of an NF from Bayesian neural networks (BNNs) and exploiting the natural likelihood function associated with any density estimator model, such as an NF, we sample the posterior distribution of flow parameters given training data and suitable a priori assumptions. We use Hamiltonian Monte Carlo methods(HMC, M. Betancourt, 2018) such as the No-U-Turn-Sampler (NUTS, M. D. Hoffman & A. Gelman, 2011) to constrain the exact posterior distribution of flow parameters given training data, prior information, and the choice of model architecture.
@@
Using the posterior samples, we predict the emulated population as a Bayesian credible interval, rather than a deterministic density function, which is expected to contain the true population-level distribution of synthesized CBC parameters with a certain posterior probability, quantifying uncertainties in both density estimation (necessary for SBI), and sample generation (required for data augmentation and feature prediction). Using population synthesis simulations of BBHs formed through common envelope (CE) evolution of isolated stellar binaries (S. S. Bavera et al., 2020), we show that the inferred credible intervals correctly quantify epistemic and aleatoric uncertainties in flow predictions by computing calibration curves, and demonstrate that they outperform an ensemble of flows trained on alternate realizations of sparse (thinned) datasets or on the same dataset with different parameter initializations. We further compute the Bayesian information criterion for different flow architectures to enable model comparison and thereby reduce the epistemic uncertainties that arise from sub-optimal architecture choices. We delineate how to marginalize the quantified uncertainties and perform model comparison among flow architectures for bias-free SBI and feature prediction using PopSynth simulations, as well as sketch additional applications in GW astronomy, whose implementations are left as futrue explorations.
@@
The rest of the paper is organized as follows. In Sec. 2, we highlight the main motivations for developing this framework, compare with previous studies in Sec. 3, and describe the PopSynth datasets used for our demonstrations in Sec. 4. We discuss in Sec. 5, NF emulators for PopSynth simulations and demonstrate applications for data-augmentation and SBI. In Sec. 6 our detailed implementation of the Bayesian normalizing flow, its training, and calibration. We document our code and outline the applications of our uncertainty quantification framework in Secs. 8, 7. In Sec. 9, we demonstrate the accuracy and interpretability of the quantified uncertainties for a fiducial BBH PopSynth dataset while also highlighting the role played by uncertainty quantification in data augmentation. We conclude in Sec. 10 with a summary of the main developments presented and the follow-up investigations planned for the near future.
2 Motivation: the necessity of UQ for high-fidelity simulators
@@
The scope and reliability of SBI and feature prediction are determined by the fidelity of the simulator and its underlying physical prescriptions. Compared to its earlier versions (S. S. Bavera et al., 2020), which were used to model some isolated BBH formation scenarios in the training sets of S. Colloms et al. (2025), the POSYDON framework (T. Fragos et al., 2023; J. J. Andrews et al., 2024) introduced significant advances in the treatment of binary stellar evolution, greatly expanding the astrophysical scope of SBI from GWs. POSYDON provides one of the most accurate descriptions of isolated CBC formation available in the literature by relying on extensive grids of binary evolution models that track stellar structure for both components using MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019) from zero-age main sequence to compact object formation, combined with advanced interpolation methods (T. Fragos et al., 2023). In POSYDONv2, further improvements are achieved by accounting for evolution across a cosmological range of metallicities, reverse mass transfer, spin-orbit misalignment and binary disruption due to supernova kicks, and the possibility of stellar mergers (J. J. Andrews et al., 2024), etc. A significant improvement in scalability, while retaining accuracy, is also achieved through the implementation of efficient machine learning algorithms. Altogether, POSYDONv2 is emerging as a tracktable PopSynth simulator for isolated CBC formation (and many other astrophysical systems of interest) which boasts one of the highest physics fidelities to date.
@@
Therefore, given the high computational cost of running a detailed binary simulation, POSYDONv2 can be assisted by NF emulators not only for SBI but also for feature prediction. For the latter, NFs can be used to augment the set of synthesized systems, which amounts to boosting sample statistics for a particular population, by learning from many smaller simulation sets at nearby initial condition grid points. In other words, instead of having to simulate millions of binaries for a single intitial condition of interest to properly sample under-represented regions of the CBC parameter space, many small simulation sets on a grid of nearby initial condition values (that are paralelizable as individual POSYDON runs), can be used to train an NF emulator, which can then generate a large reliable sample of binaries for the particular initial conditions of interest. This can facilitate scalable and accurate population feature-prediction for rarely synthesizable CBCs without having to simulate a prohibitively large number of systems. Relevant examples include extreme mass ratio systems with high chirp mass, and populations of systems at high ($\geq Z_{\odot}$) metallicities (K. Belczynski et al., 2010) or one that assumes large natal kick magnitudes (K. Belczynski et al., 2016), tracktable POSYDONv2 runs corresponding to which can be augmented by NF emulators to enable feature prediction with an accuracy is comparable to that of a much larger simulation. We show this later on for an existing dataset simulated with earlier versions of POSYDON, as proof of concept.
@@
In other words, PopSynth emulators trained on a diverse collection of POSYDONv2 runs promise novel and reliable astrophysical insights from growing CBC catalogs. However, generating arbitrarily many simulated binaries per initial condition grid point and using them to train emulators are both prohibitive, given the high computational cost of a single PopSynth run and that of NFs attempting to learn from billions or more datapoints, respectively. The latter might necessitate further downsampling of CBCs per population run in the SBI and/or feature-prediction training set. While NFs can facilitate data augmentation for a particular simulation grid point by learning from others (as we shall show later on), the predicted populations and density evaluations will be uncertain upon re-training with alternate realizations of down-sampled data or alternate initializations of flow parameters. The inherent randomness in POSYDON outputs due to numerical interpolation schemes used internally to speed up population synthesis, and sparsity in certain regions of CBC parameter space due to astrophysical bottlenecks, can be expected to further contribute to the variance in flow predictions for downsampled datasets. Emulating POSYDON runs for SBI and feature prediction, therefore, necessitates robust uncertainty quantification and marginalization to ensure reliable astrophysical explorations.
3 Previous Work
The implementation of Bayesian normalizing flows for conditional and unconditional density estimation, equipped with exact posterior sampling using HMC techniques that is presented here, and their application in the context of PoPSynth emulation, for SBI and data augmentation, and model comparison among flow architectures, is novel to the best of our knowledge. However, previous studies have developed UQ methods in an attempt to mitigate against the biases arising from uncertainties in flow-based deep learning models, which we discuss as follows.
In a previous study, C. Plunkett et al. (2025) discuss the effects of flow uncertainties in a very similar problem to ours, namely a simulation-based probe of Population III stars using next generation GW detectors. They train an ensemble of flows and average over the flow parameters from each training run, which, while capable of quantifying uncertainties, are, as they note, not strictly a Bayesian representation of model uncertainties (see also, L. Berry & D. Meger, 2023, for an implementation of a Monte Carlo dropout-based ensemble network approach on NFs, ). C. Plunkett et al. (2025) argue that their method outperforms approximate Bayesian techniques such as Fisher matrix analyses and variational inference, which are susceptible to exploring only single posterior modes, leading to poor uncertainty representation. However, their method can become intractable for a dense multi-dimensional grid of simulation initial conditions, or in the case of sparse and noisy training sets.
Similarly, D. Ruhe et al. (2022) have used NFs to parametrize the population distribution of observed BBHs and inferred the population likelihood as a function of flow parameters. They have maximized said likelihood using gradient descent and explored uncertainties about the optimal flow parameters. As relevant to their study, they have used unconditional density estimators whose parameters they have inferred in the Bayesian sense from a likelihood that is specific to the problem of GW population inference. They have not sampled the posterior distribution of NF parameters given training data in the context of generic conditional and unconditional density estimation (as needed for PopSynth emulation) nor explored the calibration and data augmentation impacts of the quantified uncertainties.
On the other hand, A. Delaunoy et al. (2024) have investigated SBI with Bayesian neural networks (BNNs) for low-budget ($O(10)$) simulations. They develop a functional prior on the flow parameters, the Bayesian model-average corresponding to which is well-calibrated by construction. Using mean field variational inference to train their BNNs, they demonstrate that the quantified uncertainties are better calibrated than the use of other priors on flow parameters. Again, they do not sample the exact posterior distribution for flow parameters.
In a different study, S. Bieringer et al. (2024) attempted Bayesian uncertainty quantification (UQ) for unsupervised (unconditional) density estimation using continuous normalizing flows (C. Chen et al., 2018; W. Grathwohl et al., 2018). They have either used approximate loss functions (Y. Lipman et al., 2023) instead of the exact log-likelihood or highly fine-tuned sampling methods (S. Bieringer et al., 2023) that have been subject to limited testing, to conduct their inference. They further restrict to a few (O(50)) posterior samples for constructing calibration statistics and do not outline any model-comparison framework. Furthermore, since they restrict to only unconditional density estimators, generalizations of their method to SBI and data augmentation in the context of PopSynth emulators and its performance on test sets excluded from training remain unexplored.
Instead, the method proposed here uses more robust, stable, and rigorously tested sampling techniques based on HMC methods, exact log-likelihoods, and calibration statistics constructed from thousands of posterior samples, which ensure stability and convergence. We sample multiple HMC chains parallely on separate GPUs to avoid exploring only one of multiple posterior modes. We further show that our method outperforms uncertainty estimation using NF ensembles for sparse datasets. While we use masked autoregressive flows (G. Papamakarios et al., 2018) for the proof-of-concept demonstration presented here, this method is generalizable to continuous flows, which are expected to enable much faster convergence for HMC sampling. As we show later on, for advanced continuous flow algorithms (such as FFJORD, W. Grathwohl et al., 2018) with scalable log density evaluations, the complexity required to emulate realistic CBC PopSynth simulations is also well within the scope of modern HMC samplers. Hence, a PopSynth emulator based on Bayesian continuous flows would not necessitate highly fine-tuned samplers such as the ones proposed by S. Bieringer et al. (2023). This endeavor is left as a future exploration.
4 Population Synthesis datasets
To demonstrate the performance and accuracy of our method, we use population synthesis simulations of merging BBHs formed through the common envelope evolution of isolated stellar binaries (S. S. Bavera et al., 2020). The formation of these systems was modeled using an earlier version of the POSYDON framework, which combines various stages of binary evolution simulated by the rapid population synthesis code COSMIC (K. Breivik et al., 2020) with the detailed stellar structure calculations performed by MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019), and is described in  (T. Fragos et al., 2023). Each population in the simulated dataset is characterized by two initial conditions, namely the common envelop ejection efficiency $\alpha$ and the BH natal spin $\chi_{b}$. The simulations were carried out over a grid of initial conditions spanned by $\alpha\in\{0.2,0.5,1.0,2.0,5.0\}$ and $\chi_{b}\in\{0,0.1,0.2,0.5\}$. We chose the population corresponding to $\alpha=1.0,\chi_{b}=0.1$ to be our test set by excluding it from both training, validation, and Bayesian inference.
5 Normalizing Flow Emulators for population synthesis simulations